# Data Matters:

### v1

@@ Abstract
Accurate predictions of base stations’ traffic load are essential to mobile cellular operators and their users as they support the efficient use of network resources and sustain smart cities and roads. Traditionally, cellular network time-series have been considered for this prediction task. More recently, exogenous factors such as points of presence and other environmental knowledge have been introduced to facilitate cellular traffic forecasting. In this study, we focus on smart roads and explore road traffic measures to model the processes underlying cellular traffic generation with the goal to improve prediction performance. Comprehensive experiments demonstrate that by employing road flow and speed, in addition to cellular network metrics, cellular load prediction errors can be reduced by as much as $56.5\%.$ The code and more detailed results are available on https://github.com/nvassileva/DataMatters.
@@
data, machine learning, mobile cellular traffic, forecasting.
@@ I Introduction
Accurate mobile cellular traffic predictions are critical to mobile operators as the efficient use of network resources and delivery of services depends largely on them. Smart cities and roads are highly dependent on the availability of mobile cellular infrastructure too. Traditionally, the traffic load at a base station (BS) is predicted based on mobile cellular key performance indicators (KPIs). Recently, with the introduction of deep learning models [1], exogenous factors, in addition to cellular time-series data, are incorporated to further facilitate the forecasting process. Typically considered are points-of-presence such as commercial centers, touristic spots or sport venues. More recent works [2] broaden this contextual modeling to include more environmental factors that characterize the surroundings of each BS. In an effort to capture mobility dynamics the handover rate is harnessed in [3]. Nonetheless, data that captures population dynamics remains mainly unexplored. Our goal is to model better the cellular traffic generation process by employing information about the density of the potential sources of cellular load together with the cellular time-series data. In contrast to the majority of the works, we focus on highways instead of urban scenarios and on short-term forecasting, which is arguably more challenging than mid- and long-term forecasting yet relevant to low-latency and high-reliability services, especially those related to smart traffic and autonomous driving.
Our major contributions can be summarized in:
  • We capture the intricate dynamics of the mobile cellular traffic generation process by combining road traffic metrics that account for the potential sources of cellular communication load with cellular traffic time series data.
@@
  • We develop a methodology for generating cellular traffic volumes based on road traffic.
@@
  • We conduct comprehensive experiments for a real-world highway under a diversity of mobile cellular and road traffic conditions, which complement our first study [4].
@@
The methodology we develop for creating cellular traffic data sets based on road traffic is explained in Section II. A description of the real highway scenario used in the experiments can be found in Section III. The traffic prediction model and the learning model we use for solving it are formulated in Section IV. We discuss the experimental setting and performance results in Section V, and conclude the paper with a summary of the main results and future prospects in Section VI.
@@ II Methodology
Measuring and recording a large variety of different KPIs is intrinsic to any communication network because the measurements are used for monitoring, operating, optimizing and maintaining the network. However, mobile operators rarely share such data because of privacy concerns. One of the very few open access data sets [5] is the Telecom Italia Big Data Challenge data [6]. Therefore, the majority of the studies, which develop machine learning methods for predicting mobile cellular traffic, routinely use it. Commonly, their goal is to predict the number of calls as a measure of the cellular network load. We follow the same approach of using calls.
@@ II-B1 Vehicular flow arrival and departure
The loop detectors of the California transportation system measure all variables in real time yet the measurements are recorded in PeMS as aggregates for the number of vehicles and as averages for the speed per five-minute intervals [^1] The five-minute interval is typical for most transportation systems. as mentioned earlier. Vehicular arrival and departures times at a BS are needed to simulate the call cellular load. We use the conventional Poisson assumption to emulate vehicular arrival times. Some more recent studies [8] validate this hypothesis with real data from highways for sparse—not congested nor experiencing bursts—traffic. Since for the large majority of the data we employ, the speed is around its maximum allowed value (70 mph), suggesting congestion is not habitual, we adopt this assumption too: we emulate the vehicles’ arrival time by exponential inter-arrival times within each five-minute interval. The vehicle’s departure time is determined by its arrival and dwell times.
@@ II-B2 Dwell time
The time spent by a vehicle in a segment of a highway served by a BS, is modeled by the range of the BS and the measured PeMS average speed. We add Gaussian noise with mean zero and standard deviation 0.05 to emulate different speeds.
@@ II-B3 Calls
New call arrivals at a BS follow a Poisson process (see [9] for an overview). Therefore, we model the calls placed by each vehicle by exponential call inter-arrival times. Each vehicle generates calls with a given rate $\lambda$, which is fixed and common for all vehicles. The call arrival process at a BS then is a non-homogeneous Poisson process, with arrival rate given by the sum of the time-varying vehicular arrival rate (different per five-minute interval) and the fixed call arrival rate.
Figure 1: A highway with merging arterial roads and BS coverage.
@@ II-B5 Flow between adjacent cells
The vehicular flows in two adjacent BSs can differ from each other. The recorded number of vehicles in one segment of the highway can be smaller or higher than that of the subsequent segment of the highway. This can be due to arterial roads that merge into the highway, see a schematic representation of such a case in Fig. 1. When the incoming flow in the target cell (BS2 in Fig. 1) is smaller than the outgoing flow from the preceding cell (BS1), we randomly remove vehicles together with their corresponding handover calls from the generated data set. The goal is to reflect the route deviation of the vehicles, namely those that leave the highway, into the network statistics too. When the flow in the target cell is larger than the adjacent, preceding cell, no adjustments are made in the handover statistics generated in the preceding cell. The generation of new calls per each vehicle arriving at the cell is done as explained earlier.
@@ III A highway scenario
Figure 2: A map of the section of US50-E El Dorado County highway used in the study (in gray). The red dots indicate the approximate location of the PeMS detectors on that segment, from Pollock Pines to Lake Tahoe Airport. The detectors are listed in Table I together with the distance between consecutive ones.
We consider a sector of the US50-E highway in El Dorado County, California; specifically, the one shown in Figure 2. The Caltrans detectors, whose measurements are used in the study, are listed in Table I in sequential order from left to right when looking at the map. Their approximate location is marked with red dots in Figure 2. The distance between each two consecutive detectors is indicated in Table I.
The density of the PeMS detectors is high at the fragments of the highway that cross populated areas (such as villages and towns; Pollac Pines and Mevers in this particular scenario) and is much lower on the highway sectors of mountainous regions. Long-range BS typically cover rural areas with sparse population, whereas short-range BSs provide coverage in densely populated (urban) regions. Since the density of the road detectors seems to be guided by similar principles—long-term density of the vehicular traffic—as the distribution of BS— population density and the particularities of the terrain—we choose the range of the BSs to match the distance between sensors. This provides us with a variety of different use cases as the BS range (under otherwise equal conditions) defines the handover rates and cell load. For simplicity, we make the assumption that the vehicular traffic is unidirectional, flowing from Pollac Pines to South Lake Tahoe.
TABLE I: PeMS US50-E Detectors Used in the Study
and the Distance Between Consecutive Ones
BS# PeMS detectors miles
1 Mainline VDS 3086071–50EB JWO Sly Park EB $7.67$
2 Mainline VDS 3086081–50EB at Riverton Barn CCTV $1.62$
3 Mainline VDS 320287–50EB At Ice House $13.27$
4 Mainline VDS 320280–Wrights Lake Rd $13.41$
5 Mainline VDS 317706–Echo Summit $3.89$
6 Mainline VDS 3054051–50EB into Luther 50/89 R. $0.99$
7 Mainline VDS 3410061–50EB JEO Pioneer Trl $3.22$
8 Mainline VDS 317715–F St 1.46
@@ IV Mobile Cellular Traffic Prediction: Definition and Solution
Our goal is to predict the mobile cellular traffic load at each BS based on the historical time-series of their mobile cellular volumes and the road traffic time-series observed in the highway segments covered by the BSs. We focus on a single machine learning model, individually trained (with the BS’s own data) for each BS along the highway since our goal is to validate the usefulness of the concept rather than develop a new method. A basic machine learning model allows to abstract the impact of the model from that of the data set and hence, it is easier to validate.
@@ IV-A Problem Formulation
Mathematically, we denote by $\mathbf{x}^{\tau}\in\mathcal{R}^{p}$ a random variable comprising historical measurements of $p$ metrics relevant to a given cell: mobile traffic volume, and/or road traffic volume and speed measurements. Given a time sequence of $M$ such historical observations $\{\mathbf{x}^{\tau-M+1},\dots,\mathbf{x}^{\tau}\}$, our objective is to learn a prediction model $\mathcal{F}$ that can forecast the future mobile traffic load $\hat{x}$ in the cell during the next time step:
@@
IV-B A basic machine learning model
Mobile cellular traffic volumes at BSs experience spatial-temporal dependencies. Since we focus on a short-term forecasting problem at a BS (1), the impact of spatial correlations is confined to the adjacent BSs. We account for this phenomenon by the incoming vehicular flow and speed.
The long short-term memory (LSTM) model [13] has been applied in many time-series prediction tasks including mobile cellular traffic volume forecasting. The model we construct is composed by an LSTM layer, followed by a fully connected feedforward neural network (a multilayer perceptron). This dense layer consists of a single unit and no activation function. It is a linear transformation that maps the final LSTM state into a single target value—the estimated number of calls in a BS.
Mathematically, the model can be represented as two functions $f(g),$ where $g(\cdot)$ is the LSTM learning structure, which transforms the input data into new features. The representation function $f(\cdot)$ maps the learned features into a cellular traffic consumption prediction for the BS under consideration.
@@ V-A Setting
V-A1 Input variables values
Real road traffic data from Caltrans detectors of the US50-E highway in El Dorado County, California are used in the experiments. The considered time period is 24-week-long from March 28, 2022 to September 9, 2022 (week 13 to week 24 of 2022). The data comprise the flow and speed from Monday to Friday of each week. There are 288 data points per day due to the 5-minute granularity.
We generate the network statistics following the methodology developed in Section II and input values as follows. New calls are generated with arrival rate $\lambda=1/5,$ one call per 5-minute interval on average. When defining the mean call duration and its variance, we are guided by [12] and [10]. We set the mean call duration to 1 minute for exponentially distributed calls and to 1 and 10 minute for the mixture of two log-normally distributed calls, each with equal weight. The variance of the log-normal calls is 3 times larger than their mean [12]. The range of the BSs is listed in Table I.
@@ V-A2 Data
All BSs experience mostly typical for their highway segment mobile volumes yet these traffic loads are quickly varying in short-term scales and are weekly different. For several BSs there are outlying from the rest weeks.
The road metrics—flow and speed—are inversely associated. In addition, their correlation varies over time.
Flow and the generated number of calls are strongly correlated. This reciprocal association, corresponds to recent [14] and earlier studies [15], which show that there exists a strong correlation between population density and mobile cellular use in urban scenarios. Research on population mobility relies on such a high correlation when using call detail records (CDRs) from mobile operators to understand and model human mobility in urban environments [16]. Furthermore, people’s density is often estimated from mobile phone network data [16], [17], [18]. Under the assumption of connected vehicles and a highway scenario—the setting of our study—the load on the BSs that serve the highway is exclusively generated by the road traffic on that highway (the applications used by the autonomous driving vehicles and their passengers). Naturally, the highest correlation is between the flow of vehicles and the new calls they place on their serving BS and correspondingly, the total number of calls. The linear association between the flow and number of handovers is often of a smaller magnitude because not all vehicles have active calls when crossing the borders between BSs, namely when arriving to a new cell.
The number of calls is inversely related to the speed as whenever the road is congested, the dwell time of the vehicle on the road is increased and consequently the probability of initiating a new call grows because we model the call arrival rate as a constant random variable (a variable with variance zero). This is in line with the common observation that the load on the mobile cellular network increases during traffic jams, although we do not model this phenomenon explicitly.

### v2

@@ Abstract
Accurate predictions of base stations’ traffic load are essential to mobile cellular operators and their users as they support the efficient use of network resources and allow delivery of services that sustain smart cities and roads. Traditionally, cellular network time-series have been considered for this prediction task. More recently, exogenous factors such as points of interest and other environmental knowledge have been explored too. In contrast to incorporating external factors, we propose to learn the processes underlying cellular load generation by employing population dynamics data. In this study, we focus on smart roads and use road traffic measures to improve prediction accuracy. Comprehensive experiments demonstrate that by employing road flow and speed, in addition to cellular network metrics, base station load prediction errors can be substantially reduced, by as much as $56.5\%.$ The code, visualizations and extensive results are available on https://github.com/nvassileva/DataMatters.
@@
data, machine learning, mobile cellular traffic, forecasting, population dynamics.
@@ I Introduction
Accurate predictions of base stations’ (BSs) traffic load are essential to mobile cellular operators and their users as they support the efficient use of network resources and allow delivery of services that sustain smart cities and roads. Significant progress has been made in developing deep learning architectures for cellular traffic predictions [1]. Nonetheless, learning useful representations is generally based on mobile cellular key performance indicators (KPIs) much like it has been done in the past with classic statistical models for time-series predictions. Recent works incorporate exogenous information namely, auxiliary data about extrinsic factors that might have an effect on the cellular load [2]. Typically, these are points of interest (locations with high user activity). In some of the latest contributions [3], [4], the contextual description is enriched by including spatial, temporal and semantic relationships that describe the BSs and their surroundings.
TABLE I: Employing Population Dynamics in Addition to Cellular Data Improves Prediction Performance by at Least $5.7\%$ and by More Than $50\%$
BS MAE MSE MAPE RMSE
# min mdn max min mdn max min mdn max min mdn max
3054051 16.0 15.0 14.0 29.7 29.4 29.6 27.4 23.6 27.4 16.1 16.0 16.1
3086071 33.9 34.0 34.7 56.3 56.5 56.6 38.2 38.4 40.5 33.9 34.0 34.1
3086081 13.9 14.1 14.4 25.1 25.2 25.4 12.5 14.1 16.3 13.5 13.5 13.6
317706 16.6 16.4 15.0 33.8 33.2 33.5 15.8 13.7 18.7 18.6 18.3 18.4
317715 8.0 8.4 7.8 14.0 13.3 11.1 15.6 15.4 20.8 7.3 6.9 5.7
320280 18.3 18.4 27.4 35.3 36.8 40.0 12.5 12.5 14.3 19.6 20.5 22.5
320287 23.0 21.4 19.5 38.9 36.8 35.1 25.5 23.5 23.4 21.8 20.5 19.4
3410061 12.5 12.0 12.0 20.6 20.9 20.6 30.4 31.9 33.6 10.9 11.0 10.9
aMean Absolute (Percentage) Error (MA(P)E), (Root) Mean Square Error ((R)MSE).
bThe three values listed per error type are the minimum, median and maximum.
In contrast to accounting for external factors, we propose to learn the intrinsic forces that govern the mobile cellular traffic generation. We hypothesize that by learning the processes underlying BS load generation, the accuracy with which cellular volume is predicted can be scaled up significantly. To this end we propose to employ data that characterizes population dynamics namely, the potential sources of cellular load and their fluctuations over time in addition to cellular KPIs.
We focus on highways and teleoperated and autonomous driving; hence, on short-term forecasting, which is arguably more challenging than mid- and long-term but relevant to low-latency and high-reliability services, especially those related to smart traffic and vehicles. We combine road data with mobile cellular KPIs to enrich the cellular time-series with population dynamics data. The computational cost of using road statistics is negligible, especially compared to recent studies [3], which incorporate a large knowledge graph. The combined improvement attained from using this graph and the developed deep learning architecture in [3] when contrasted with state-of-the-art approaches is at most $18\%$ [3]. In our study, we record improvements above $50\%$, Table I, due to the use of road statistics alone. The communication overhead between BSs and roadside units, which are typically used in smart roads where real-time traffic statistics are within reach, is of little relevance.
Our primary contributions are:
  • We are the first to propose and capture the phenomena intrinsic to mobile cellular traffic generation. We employ road traffic metrics to gauge the potential sources of BS load in highways together with cellular time-series. This leads to substantial improvements in prediction accuracy.
@@
  • We develop a methodology for generating cellular traffic volumes in highways based on real-life road traffic data.
@@
  • We conduct comprehensive experiments for a real-world highway under quickly fluctuating road traffic load, seasonal changes and a diversity of mobile communication volumes, which complement our first study [5].
  • We examine highways and address short-term mobile cellular traffic predictions in contrast to the extensively studied urban scenario and long-term cellular forecasting.
@@
The methodology we develop for creating cellular traffic data sets based on road traffic is explained in Section II. A description of the real highway scenario used in the experiments can be found in Section III. The traffic prediction problem and the implemented learning model are formulated in Section IV. We discuss the experimental setting, methodology and protocol, and performance results in Section V. The novelty and impact of the approach we propose are highlighted in Section VI. The main results and future prospects are summarized in Section VII. The code, extensive results and visualizations are available on https://github.com/nvassileva/DataMatters.
@@ II Methodology
Measuring and recording a large variety of different KPIs is intrinsic to any communication network because the measurements are used for monitoring, operating, optimizing and maintaining the network. However, mobile operators rarely share such data because of privacy concerns. One of the very few open access data sets [6] is the Telecom Italia Big Data Challenge [7]. Therefore, the majority of the studies, which develop machine learning methods for predicting mobile cellular traffic, routinely use it [2]. Commonly, their goal is to predict the number of calls as a measure of the cellular network load. We adopt the same approach of using the total calls number as BS load indicator too—we expect the general trends we observe in this study to remain the same independent of the load traffic measure used.
@@ II-B1 Vehicular flow arrival and departure
The loop detectors of the Caltrans system measure all variables in real time yet the measurements are recorded in PeMS as aggregates for the number of vehicles and as averages for the speed per five-minute intervals, as mentioned earlier. Vehicular arrival and departures times at a BS are needed to simulate the call cellular load. The classic Poisson assumption about arrival times has been validated in [9] with real data from highways for sparse—not congested nor experiencing bursts—traffic. Since for the large majority of the data we employ, the speed is around its maximum allowed value, suggesting congestion is not habitual, we adopt this assumption too. In effect, we emulate each vehicle’s arrival time by exponential inter-arrival times within each five-minute interval. The vehicle’s departure time is determined by its arrival and cell dwell times.
@@ II-B2 Dwell time
The time spent by a vehicle in a segment of a highway served by a BS, is modeled by the BS’s range and the measured PeMS average speed. We add Gaussian noise with mean zero and standard deviation 0.05 to emulate different speeds.
@@ II-B3 Calls
New call arrivals at a BS follow a Poisson process as demonstrated empirically, analytically and by simulation [10]. Therefore, we model the calls placed by each vehicle by exponential call inter-arrival times. Each vehicle generates calls with a given rate $\lambda$, common for all vehicles. Although $\lambda$ is fixed, the new call arrival process at a BS is a non-homogeneous Poisson process, with time-varying arrival rate given by the sum of the time-varying vehicular arrival rate (different per five-minute interval) and the fixed call arrival rate. As this is a random process, each passing vehicle might generate no service request, at most one, or multiple requests while driving through the segment of the freeway covered by the BS. Furthermore, during congestion periods, the simulated load on the network grows correspondingly (see Section V-A2).
@@
Figure 1: A map of the section of US50-E El Dorado County freeway used in the study (in gray). The red dots indicate the approximate location of the PeMS detectors on that segment, from Pollock Pines to Lake Tahoe Airport. A larger image is available on https://github.com/nvassileva/DataMatters.
@@ II-B5 Flow between adjacent cells
The vehicular flows in two adjacent BSs can differ from each other. The recorded number of vehicles in one segment of the highway can be smaller or higher than that of the subsequent segment. This can be due to arterial roads that merge into the highway. When the incoming flow in the target cell is smaller than the outgoing flow from the preceding cell, we randomly remove vehicles together with their corresponding handover calls from the generated data set. The goal is to reflect the route deviation of the vehicles, namely those that leave the highway, into the network statistics too. When the flow in the target cell is larger than the adjacent, preceding cell, no adjustments are made in the handover statistics generated in the preceding cell.
@@ III A highway scenario
We consider a sector of the US50-E freeway in El Dorado County, California; specifically, the one shown in Figure 1. The Caltrans detectors, whose measurements are used in the study, are listed in Table II in sequential order together with the distance between each two consecutive detectors.
The density of the PeMS detectors is high at the fragments of the highway that cross populated areas (such as villages and towns; Pollac Pines and Mevers in this particular scenario) and is much lower on the highway sectors of mountainous regions. Long-range BS typically cover rural areas with sparse population, whereas short-range BSs provide coverage in densely populated (urban) regions. Since the density of the road detectors seems to be guided by similar principles—long-term density of the vehicular traffic—as the placement of BSs—population density and the particularities of the terrain—we choose the range of the BSs to match the distance between sensors. This provides us with a variety of diverse use cases. The road segments exhibit different capacity, average vehicular density and propensity for congestion, which conditions together with the different BSs’ ranges model dynamic, time- and location-varying call volumes. For simplicity, we make the assumption that the vehicular traffic is unidirectional, flowing from Pollac Pines to South Lake Tahoe.
TABLE II: PeMS US50-E Detectors Used in the Study
and the Distance Between Consecutive Ones.
BSs Have the Same ID as the Sensors.
PeMS detectors miles
Mainline VDS 3086071–50EB JWO Sly Park EB $7.67$
Mainline VDS 3086081–50EB at Riverton Barn CCTV $1.62$
Mainline VDS 320287–50EB At Ice House $13.27$
Mainline VDS 320280–Wrights Lake Rd $13.41$
Mainline VDS 317706–Echo Summit $3.89$
Mainline VDS 3054051–50EB into Luther 50/89 R. $0.99$
Mainline VDS 3410061–50EB JEO Pioneer Trl $3.22$
Mainline VDS 317715–F St 1.46
@@ IV Mobile Cellular Traffic Prediction: Definition and Solution
We assess the efficacy of our proposed approach by training a suitable machine learning model with historical cellular data together with versus without incorporating road traffic data. The classic machine learning model we implement is individually trained (with the BS’s own data) for each BS along the highway.
@@ IV-A Problem Formulation
Mathematically, we denote by $\mathbf{x}^{\tau}\in\mathcal{R}^{p}$ a random variable comprising historical measurements of $p$ metrics relevant to a given cell: cellular traffic volume or cellular and road traffic volume and speed measurements. Given a time sequence of $M$ such historical observations $\{\mathbf{x}^{\tau-M+1},\dots,\mathbf{x}^{\tau}\}$, our objective is to learn a prediction model $\mathcal{F}$ that can forecast the future mobile traffic load $\hat{x}$ in the cell during the next time step:
@@
IV-B A classic machine learning model
Mobile cellular traffic volumes at BSs experience spatial-temporal dependencies. Since we focus on a short-term forecasting problem at a BS (1), the impact of correlations is confined to the adjacent BSs. We account for this phenomenon by the incoming vehicular flow and speed.
Refer to captionRefer to caption
Figure 2: BS 320287 Tuesday’s and weekly fluctuations of cellular load for two specific weeks.
The classic long short-term memory (LSTM) model has been broadly applied to time-series prediction tasks either as a singular learning structure or as a basic component of more complex models. Furthermore, LSTM has been shown to excel in very short-term cellular forecasting [14], which motivates our choice.
The model we construct is composed by an LSTM layer, followed by a fully connected feedforward neural network (a multilayer perceptron). This dense layer consists of a single unit and no activation function. It is a linear transformation that maps the final LSTM state into a single target value—the estimated number of calls in a BS.
Mathematically, the model can be represented as two functions $f(g),$ where $g(\cdot)$ is the LSTM learning structure, which transforms the input data into new features. The representation function $f(\cdot)$ maps the learned features into a cellular traffic load prediction for the BS under consideration.
@@ V-A Setting
V-A1 Input variables’ values
Real road traffic data from Caltrans detectors of the US50-E highway in El Dorado County, California are used in the experiments. The considered time period is 24-week-long, from March 28, 2022 to September 9, 2022 (week 13 to week 36 of 2022). The data comprise the flow and speed from Monday to Friday of each week. There are 288 data points per day due to the 5-minute granularity.
We generate the network statistics following the methodology developed in Section II and input values as follows. New calls are generated with arrival rate $\lambda=1/5,$ one call per 5-minute interval on average. When defining the mean call duration and its variance, we are guided by [13] and [11]. We set the mean call duration to 1 and 10 minutes for a mixture of two log-normally distributed calls, each with equal weight. The variance of the log-normal calls is 3 times larger than their mean [13]. The range of the BSs is listed in Table II.
@@ V-A2 Data
All BSs experience a highly dynamic road and cellular traffic volumes. The period from March to September accounts for different seasons and correspondingly, diverse road traffic patterns that reflect seasonal changes and holiday periods. As a result, each BS has different daily and weekly profiles along the studied period. Furthermore, for all BSs the flow and communication load vary quickly within the 5-minute intervals. Figure 2 shows the profile of call load 1) for the same location and day of a week but different weeks and 2) for all work days of a week and different weeks for a BS as an example.
The correlation between the road metrics—flow and speed—varies greatly between segments of the highway and with time: strongly inverse, reciprocal or lacking linear association. This variability can be explained by the flow-speed correlation dependence on the capacity of the road segment, vehicular flow, terrain, location and time.
Flow and the generated number of calls are strongly correlated. This reciprocal linear association, corresponds to recent [15] and earlier studies [16], which show that there exists a strong correlation between population density and mobile cellular use in urban scenarios. Research on population mobility, in effect, relies on such a high correlation when using call detail records from mobile operators to understand and model human mobility in urban environments [17]. Furthermore, people’s density is often estimated from mobile phone network data [17], [18]. Under the assumption of connected vehicles and a highway scenario—the setting of our study—the load on the BSs that serve the highway is exclusively generated by the road traffic on that highway (services used by the autonomous driving vehicles and their passengers). Naturally, the highest correlation is between the flow of vehicles and the new calls they place on their serving BS and correspondingly, the total number of calls. The correlation between the flow and number of handovers is often of a smaller magnitude as not all vehicles have active calls when crossing the borders between BSs, namely when arriving to a new cell.
The number of calls is in general inversely related to the speed as whenever the road is congested, the dwell time of the vehicle on the road is increased and consequently the probability of initiating a new call grows. This is in line with the common observation that the load on the mobile cellular network increases during traffic jams.